# Probing the Origins I: Generalised Additive Model inference of birth radii for Milky Way stars in the solar vicinitythanks: Catalogue reference

### v1

@@ 2.3.2 The general form of a GAM
$g(\mu)=\beta_{0}+s_{1}(x_{1})+s_{2}(x_{2})+\ldots+s_{p}(x_{p})+\epsilon,$ (1)
in which $g$ is a specified link function that relates the linear predictor to the expected value of the response variable, $\mu$ represents the expected value of the response variable, $\beta_{0}$ is the intercept term, $s_{i}$ represents the smooth function associated with the $i$th predictor variable, and $x_{i}$ denotes the $i$th predictor variable.
@@
2.3.3 A GAM applied to radial metallicity profiles
@@
We employed a GAM to extend existing theoretical models that map the radial metallicity profile of the Galaxy (such as the ones presented by Magrini et al., 2009), specifically considering the Galactocentric distance ($R$) as the response variable. Our objective was to accommodate stars of diverse ages and evaluate whether they currently reside at their birth radii. To achieve this, we utilised the mgcv package in r to build the GAM, which is formulated as:
@@
In Appendix B, we provide evidence to support our decision to use [Fe/H] as the primary chemical feature in our GAM model. We achieve this by presenting an analysis of the GAM model that exclusively considers [Mg/H], thereby isolating its effects from those of Fe. Additionally, the appendix includes a comparative study between two GAM models: one using [Fe/H] and the other [Mg/H]. This comparison aims to show how the choice of chemical feature impacts the estimated Galactocentric birth distances of stars. The main body of our paper focuses on the model described in Eq. 3, which uses [Fe/H].
@@
Figure 6: New metallicity profile grids generated using the GAM. Each curve illustrates the radial metallicity profile of stars formed at different epochs in the Universe’s history, ranging from 0 to $t$ Gyr (age of the Universe). It is also worth mentioning that each subplot has its $y$-axis ($R$) truncated up to 25 kpc. Also, it is worth mentioning that the GAM does not limit the radii it can reach in a regression, therefore we chose to display up to physically meaningful radii.
Figure 6 illustrates the radial metallicity profiles estimated from the GAM across a more dense grid of ages spaced at a 1 Gyr interval. For the $y$-axis, even though the GAM is not restricted by physical limitations, we display the response variable $R$ within the meaningful Galactocentric range of 0 to 25 kpc. This value encompasses the upper estimates for the MW’s Galactocentric radius, which is generally thought to range from 17 to 25 kpc; see e.g. Huang et al. (2016) for an estimated radius based on the rotation curve of the MW, and Lian et al. (2024) for a smaller estimation of 17 kpc. Additionally, the open clusters used to create the models described by Magrini et al. (2009), reach up to 22 kpc. Thus, our choice of 25 kpc is a safe choice, encompassing even the largest estimates for the MW.
@@
2.3.4 A GAM applied to Gaia-ESO sample of stars
The model described in Eq. 3, with functions fitted to the Magrini et al. (2009) models, was then applied to the 1460 stars in our sample (described in Sect. 2.1). The regression is performed directly using the stars’ parameters: [Fe/H] and $t$, where $t$ is the age of the Universe when the star was formed, calculated as follows.
$t=t_{\text{u}}-\overline{t}_{\star},$ (4)
where $t_{\text{u}}$ is the current estimated age of the Universe (adopted value: 13.8 Gyr; see Planck Collaboration et al., 2020); and $\overline{t}_{\star}$ is the median age of the star estimated using unidam (Mints & Hekker, 2017, 2018), as described in Sect. 2 and, in more depth, by Dantas et al. (2023).
The GAM is then useful not only for providing finer grids – as described previously – but also for directly estimating the most likely birth radius ($R_{\rm b}$) for a star with these parameters. To estimate the uncertainties for $R_{\rm b}$, we applied the GAM after using a bootstrap method, resampling each star’s parameters 1000 times. The parameters used for the bootstrap were [Fe/H] and $t$, along with their respective uncertainties. The bootstrapping procedure closely follows the methodology outlined in Sect. 2.2 of Dantas et al. (2023) and therefore we omit the detailed steps here for simplicity. This approach not only provides $R_{\rm b}$ for each star, but also yields the associated quantiles, standard deviation, and other statistical measures.
Therefore, whenever we use $R_{\rm b}$, we refer to the birth radius estimation derived from the GAM. Meanwhile, $\langle R_{\rm{b}}\rangle$ will denote the median birth radius for each star, as estimated through the bootstrapping method and the GAM.
@@ 3 Analysis and results
This Section begins with an examination of the chemo-dynamic properties of our sample. This analysis is crucial as it helps to identify stars that may not belong to the Galactic thin disc, and therefore cannot have their birth radii estimated using our selected chemical evolution models. This analysis is performed in Sect. 3.1 and we briefly describe the final sample in Sect. 3.1.3, after the removal of halo and thick disc intruders.
@@ 3.1.1 Distinguishing between disc and halo stars
These outliers exhibit consistent characteristics across all plots. Notably, as metallicity decreases, a proportionally higher number of outliers are identified. Stars in groups 4 and 3, in particular, are the most metal-poor and have the largest median ages. In addition to their low metallicities and high ages, some of the outliers in these groups possess $\langle L_{z}\rangle$ and $\langle E_{t}\rangle$ values consistent with remnants of accretion and/or merger events. These characteristics align with expectations from the Gaia-Enceladus merger (GE; Belokurov et al. 2018; Helmi et al. 2018; and see Giribaldi & Smiljanic 2023 for a precise time estimate of the GE merger, suggesting that it was completed $9.6\pm 0.2$ Gyr ago). Since the focus of this paper is not to explore the various merger or accretion scenarios that have influenced MW’s evolution, we will not include these outlier stars in our analysis.
@@
Figure 8: Lindblad diagram for all the six groups retrieved from the HC, ordered by descending median metallicity $\langle\rm{[Fe/H]}\rangle$. A vertical dashed line is depicted at $\langle L_{z}\rangle$=0, which denotes the stars with retrograde orbits around the Galactic centre. Stars with $\langle L_{z}\rangle$ between 1 and 3 show stars that are probably well constrained within the MW’s disc, whereas those with 0 $\leq$ $\langle L_{z}\rangle$$<1$ have a higher probability of not being part of the disc, and those with $\langle L_{z}\rangle$$<0$ probably come from the MW’s halo. A uniform range is maintained on both the $x$ and $y$-axes, serving to direct the reader’s attention towards outliers manifesting within lower metallicities. The colour scheme remains consistent with the preceding figures.
@@ 3.1.2 Assessing thin and thick disc membership
Table 2: Table expressing the number of stars classified in the thin or thick disc according to the prescription adopted here for each metallicity-stratified group (“G”).
@@ 3.1.3 The final sample post chemo-dynamic assessment
It is noteworthy that, after excluding stars potentially belonging to the Galactic halo and those that seem to belong to the thick disc, the median age, $\overline{t}_{\star}$, for the groups changes, and the variation becomes smaller. The difference between the youngest and oldest groups is reduced to $\sim$ 2 Gyr, which is within the expected age uncertainty ($\approx$ 1-2 Gyr; see Mints & Hekker, 2018). These are median values, and we will discuss other effects of age and kinematics in Sect. 3.3. Additionally, Group 4, which is the second most metal-poor, is now the oldest when comparing their age distributions. We display both the Kiel diagram for entire final sample and the distribution of these ages in Figs. 10 and 11, respectively. The tracks shown in Fig. 10 are PARSEC isochrones retrieved from the CMD web interface [^9] http://stev.oapd.inaf.it/cgi-bin/cmd. Regarding Fig. 11, please note that this distribution is slightly different from the one in the adjacent upper axis of Fig. 2, which contains the full sample, not only the stars in the Galactic thin disc.
@@
Figure 11: The age distribution for the stars belonging to the Galactic disc is illustrated in the shape of Gaussian kernel densities for easier view (instead of histograms). The $x$-axis has been truncated to represent a physically meaningful range of stellar ages. The vertical dashed lines in corresponding colours show the computed median ages for each subgroup in our disc stellar sample. The colour scheme is consistent with the preceding figures.
3.2 Birth radii estimation vs. current Galactocentric distances
@@
Figure 12: Estimated Galactocentric birth radii, $\langle R_{\rm{b}}\rangle$, in contrast to the present guiding radii, $\langle R_{\rm{g}}\rangle$, both measured in kpc, encompassing all stars in our sample that are part of the disc, according to Fig. 7. Each distinct subplot delineates a specific star group, classified using Hierarchical Clustering (as illustrated in Fig. 1). The subplots are arranged in descending order of the median iron abundance, $\langle\rm{[Fe/H]}\rangle$. At the upper part of each subplot, $\langle\rm{[Fe/H]}\rangle$ and $\overline{t}_{\star}$ are displayed. Furthermore, the overall standard deviations ($\sigma$) for $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$ are exhibited at the lower right corner of each subplot. We overlay the 2D-Gaussian kernel densities to illustrate the concentration patterns of the stars in each subplot. A reference black dashed line is incorporated, signifying the point of parity between both radii ($\langle R_{\rm{b}}\rangle$=$\langle R_{\rm{g}}\rangle$). The colour scheme remains consistent with the preceding figures.
@@
3.2.1 Challenges in predicting birth radii: limitations of the chemical enrichment models and the GAM
Upon examining the groups depicted in Fig. 12, particularly those with sub-solar metallicities (i.e., groups 5, 4, and 3), we observe that some stars exhibit relatively large birth radii ($\langle R_{\rm{b}}\rangle$ $\sim$ 28 kpc), exceeding the estimated radius of the MW, which is approximately 25 kpc (as previously discussed in Sect. 2.3.3). One possible explanation is contamination from thick disc stars, which poses challenges for elimination without significantly affecting the completeness of the thin disc population. However, if these stars predominantly belong to the thin disc, as seems more likely for the majority, alternative explanations may apply. Rather than interpreting this as evidence of stars originating outside the Galactic disc, it more likely reflects limitations in the chemical evolution models’ ability to accurately account for the age-metallicity relationship in certain stars. Estimating accurate stellar ages, particularly for field stars, remains a significant challenge. We must also consider potential mismatches between model predictions and the actual metallicity evolution in the outer regions of the Galactic thin disc. These combined limitations lead to predictions that some stars have birth radii larger than the known extent of the MW. While the precise birth radii for these stars may be inaccurate, the broader conclusion – that they formed in the outer disc and migrated inward – remains robust.
@@ 3.2.2 Galactocentric migration patterns across different stellar groups
Figure 12 shows that $\overline{t}_{\star}$ decreases as $\langle\rm{[Fe/H]}\rangle$ increases (except for Group 3), accompanied by a reduction in the overall variance of both $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$. In other words, as we look at groups with larger metallicities, the stars seem to cluster and become more tightly grouped within a narrower range of both $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$ (note in particular that while the range of $\langle R_{\rm{g}}\rangle$ values is the same in all panels, the range of $\langle R_{\rm{b}}\rangle$ values increases from the top left to the bottom right). The clustering in $\langle R_{\rm{b}}\rangle$ for the stars with high metallicity can be easily understood with a glance at the gradients displayed in Figs. 5 and 6. Old stars of high metallicity can only form in the inner regions of the Galaxy (as we discussed previously in Dantas et al., 2023). In addition, we note that a decrease in the number density of metal-rich stars towards the outer disc can be seen in the sample of other surveys as well (e.g. Hayden et al., 2015; Gaia Collaboration et al., 2023).
@@ 3.2.3 Other signs of migration
Figure 13: Variation of Galactocentric distance vs. median eccentricity ($\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$  $\times$ $\langle e\rangle$) across stellar groups, arranged in descending order of median metallicity $\langle\rm{[Fe/H]}\rangle$. The range of eccentricity ($\langle e\rangle$) on the $x$-axis remains consistent across all subplots, while the $y$-axis retains an unconstrained scale, enhancing the visualisation of the pronounced $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$ dispersion. Corresponding to Fig. 12, a horizontal dashed line is positioned at $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$=0 to accentuate the region where Galactocentric distances coincide. To illuminate the 2D distribution of stars within the designated parameter space, 2D-Gaussian kernel densities are overlaid upon the scatter markers. Consistent with prior visuals, the colours in this figure maintain the same scheme.
Figure 14: Median maximum Galactic azimuthal height, $\langle Z_{\rm{max}}\rangle$, versus the median eccentricity, $\langle e\rangle$, for all disc stars in the six groups, ordered by decreasing metallicity from top left to bottom right. 1D-Gaussian kernel density distributions for $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$ are shown adjacent to the main subplots, with medians marked by dotted lines and annotated values. The main subplots display all disc stars, with thin (300 pc) and thick disc (900 pc) scale heights (McMillan, 2017) marked by dashed and dot-dashed lines, respectively, and a 2D Gaussian kernel density plot indicating higher density regions. Consistent $x$- and $y$-axis ranges highlight the shift in distributions across metallicity groups, with the proportion of stars with larger $\langle Z_{\rm{max}}\rangle$ increasing as metallicity decreases.
@@
Therefore, we analyse our stars’ $\langle e\rangle$ and $\langle Z_{\rm{max}}\rangle$ to explore whether these parameters are connected to migration patterns. Figure 13 presents the variations in radii, $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$ (or simply, $\Delta R$), against $\langle e\rangle$ for all groups in our sample. This figure provides alternative an view of the results shown in Fig. 12 by incorporating $\langle e\rangle$. In Fig. 13, the eccentricities of stars across all groups appear constrained to $\sim$0.4, though Group 6 contains a few outliers with higher eccentricities. As we move toward metal-poor groups, the variation in $\Delta R$ increases, but $\langle e\rangle$ remains largely unchanged.
@@
Another way to visualise the behaviour of the stars in our sample is by plotting $\langle Z_{\rm{max}}\rangle$ versus $\langle e\rangle$, which is displayed in Fig. 14. The subplots in Fig. 14 are organised by decreasing metallicity, from top left to bottom right. The scale heights of the thin and thick discs are indicated in the figure by dashed and dot-dashed lines at 0.3 and 0.9 kpc, respectively (estimates from McMillan, 2017). The consistent axis ranges across all subplots allow us to observe the shifting distributions as metallicity decreases, notably the increasing proportion of stars with larger $\langle Z_{\rm{max}}\rangle$ values as metallicity decreases, especially for Groups 4 and 3. In fact, the median $\langle Z_{\rm{max}}\rangle$ for Group 4 is approximately at the scale height of the thick disc, and significantly above it for Group 5. The standard deviations for both $\langle e\rangle$ and $\langle Z_{\rm{max}}\rangle$ are displayed in each subplot, demonstrating that the variance in $\langle Z_{\rm{max}}\rangle$ increases as metallicity decreases, while the variance in $\langle e\rangle$ remains pretty constant. As for $\langle e\rangle$, it is noticeable that the medians are all $<0.2$, suggesting that most stars have near-circular orbits, but all have somewhat low $\langle e\rangle$, almost all having $\langle e\rangle$ $<0.4$; few outliers can be seen in Groups 2 and 6.
@@
As metallicity decreases, a greater number of stars appear to reach larger $\langle Z_{\rm{max}}\rangle$ values. This could be attributed to residual interlopers from the thick disc not previously identified in Sect. 3.1.2. However, if most of these stars indeed belong to the Galactic thin disc, as expected, this trend could be potentially explained by two factors: (i) stars in the most metal-poor groups tend to concentrate at older ages (as will be discussed in Sect. 3.3.1 and detailed in Table 10), giving them more time to interact with Galactic structures (e.g. giant molecular clouds), which may contribute to disc thickening, as shown by Roškar et al. (2013) – however, this applies mainly to stars churned outward, and, as we discuss in Sect. 3.3, the proportion of such stars in these metal-poor groups is low; (ii) the interaction between the outer disc and the first pericentric passage of Sagittarius, which occurred around 6 Gyr ago. According to Das et al. (2024), stars older than 6 Gyr are significantly kinematically hotter than their younger counterparts. Indeed, our observations indicate that the stars in the most metal-poor groups – those likely formed at greater Galactocentric distances – have higher $\langle Z_{\rm{max}}\rangle$ values. This is consistent with the hypothesis that stars formed at larger radii are more susceptible to the influence of interactions with Sagittarius, leading to dynamic heating. Furthermore, as we show and discuss in Sect. 3.3, a substantial proportion of metal-poor stars appear to be migrating inward from the outskirts of the MW.
@@
To further explore this hypothesis, we depict $\langle Z_{\rm{max}}\rangle$ vs. $\overline{t}_{\star}$ in Fig. 15, stratified by the HC groups. It is evident in Fig. 15 that stars aged 6 Gyr or older exhibit a higher proportion of larger $\langle Z_{\rm{max}}\rangle$ values, especially in the more metal-poor groups, such as Groups 5, 4, and 3. Overall, if the majority of these stars are not interlopers from the thick disc, we propose that this increased variation in $\langle Z_{\rm{max}}\rangle$ is likely due to a combination of outward churning and the dynamic influence of the pericentric passage of Sagittarius.

### v2

@@ 2.3.2 The general form of a generalised additive model
$g(\mu)=\beta_{0}+s_{1}(x_{1})+s_{2}(x_{2})+\ldots+s_{p}(x_{p})+\epsilon.$ (1)
Here, $g$ is a specified link function that relates the linear predictor to the expected value of the response variable, $\mu$ represents the expected value of the response variable, $\beta_{0}$ is the intercept term, $s_{i}$ represents the smooth function associated with the $i$th predictor variable, and $x_{i}$ denotes the $i$th predictor variable.
@@
2.3.3 A generalised additive model applied to radial metallicity profiles
@@
We employed a GAM to extend existing theoretical models that map the radial metallicity profile of the Galaxy (such as the ones presented by Magrini et al., 2009), specifically considering the Galactocentric distance ($R$) as the response variable. Our objective was to accommodate stars of diverse ages and evaluate whether they currently reside at their birth radii. To achieve this, we utilised the mgcv package in r to build the GAM, which is formulated as
@@
In Appendix B, we provide evidence to support our decision to use [Fe/H] as the primary chemical feature in our GAM. We achieve this by presenting an analysis of the GAM that exclusively considers [Mg/H], thereby isolating its effects from those of Fe. Additionally, the appendix includes a comparative study between two GAMs: one using [Fe/H] and the other [Mg/H]. This comparison aims to show how the choice of chemical feature impacts the estimated Galactocentric birth distances of stars. The main body of our paper focuses on the model described in Eq. 3, which uses [Fe/H].
@@
Figure 6: New metallicity profile grids generated using the GAM. Each curve illustrates the radial metallicity profile of stars formed at different epochs in the Universe’s history, ranging from 0 to $t$ Gyr (age of the Universe). It is also worth mentioning that each subplot has its $y$-axis ($R$) truncated up to 25 kpc. Also, it is worth mentioning that the GAM does not limit the radii it can reach in a regression. Therefore, we chose to display up to physically meaningful radii.
Figure 6 illustrates the radial metallicity profiles estimated from the GAM across a more dense grid of ages spaced at a 1 Gyr interval. For the $y$-axis, even though the GAM is not restricted by physical limitations, we display the response variable $R$ within the meaningful Galactocentric range of 0 to 25 kpc. This value encompasses the upper estimates for the MW’s Galactocentric radius, which is generally thought to range from 17 to 25 kpc [see e.g. Huang et al. (2016) for an estimated radius based on the rotation curve of the MW and Lian et al. (2024) for a smaller estimation of 17 kpc]. Additionally, the open clusters used to create the models described by Magrini et al. (2009), reach up to 22 kpc. Thus, our choice of 25 kpc is a safe choice, encompassing even the largest estimates for the MW.
@@
2.3.4 A generalised additive model applied to a Gaia-ESO sample of stars
The model described in Eq. 3, with functions fitted to the Magrini et al. (2009) models, was then applied to the 1460 stars in our sample (described in Sect. 2.1). The regression is performed directly using the stars’ parameters: [Fe/H] and $t$, where $t$ is the age of the Universe when the star was formed, calculated as follows:
$t=t_{\text{u}}-\overline{t}_{\star}.$ (4)
Here, $t_{\text{u}}$ is the current estimated age of the Universe (adopted value: 13.8 Gyr; see Planck Collaboration et al., 2020), and $\overline{t}_{\star}$ is the median age of the star estimated using unidam (Mints & Hekker, 2017, 2018), as described in Sect. 2 (and in more depth by Dantas et al., 2023).
The GAM is then useful not only for providing finer grids – as described previously – but also for directly estimating the most likely birth radius ($R_{\rm b}$) for a star with these parameters. To estimate the uncertainties for $R_{\rm b}$, we applied the GAM after using a bootstrap method, resampling each star’s parameters 1000 times. The parameters used for the bootstrap were [Fe/H] and $t$, along with their respective uncertainties. The bootstrapping procedure closely follows the methodology outlined in Sect. 2.2 of Dantas et al. (2023), and therefore we omit the detailed steps here for simplicity. This approach not only provides $R_{\rm b}$ for each star, but also yields the associated quantiles, standard deviation, and other statistical measures.
Therefore, whenever we use $R_{\rm b}$, we refer to the birth radius estimation derived from the GAM. Meanwhile, $\langle R_{\rm{b}}\rangle$  denotes the median birth radius for each star, as estimated through the bootstrapping method and the GAM.
@@ 3 Analysis and results
This Section begins with an examination of the chemo-dynamic properties of our sample. This analysis is crucial, as it helps identify stars that may not belong to the Galactic thin disc and therefore cannot have their birth radii estimated using our selected chemical evolution models. This analysis is performed in Sect. 3.1 and we briefly describe the final sample in Sect. 3.1.3, after the removal of halo and thick disc intruders.
@@ 3.1.1 Distinguishing between disc and halo stars
These outliers exhibit consistent characteristics across all plots. Notably, as metallicity decreases, a proportionally higher number of outliers are identified. Stars in groups 4 and 3, in particular, are the most metal-poor and have the largest median ages. In addition to their low metallicities and high ages, some of the outliers in these groups possess $\langle L_{z}\rangle$ and $\langle E_{t}\rangle$ values consistent with remnants of accretion and/or merger events. These characteristics align with expectations from the Gaia-Enceladus merger (GE; Belokurov et al. 2018; Helmi et al. 2018; and see Giribaldi & Smiljanic 2023 for a precise time estimate of the GE merger, suggesting that it was completed $9.6\pm 0.2$ Gyr ago). Since the focus of this paper is not to explore the various merger or accretion scenarios that have influenced the MW’s evolution, we did not include these outlier stars in our analysis.
@@
Figure 8: Lindblad diagram for all six groups retrieved from the HC, ordered by descending median metallicity $\langle\rm{[Fe/H]}\rangle$. A vertical dashed line is depicted at $\langle L_{z}\rangle$=0, which denotes the stars with retrograde orbits around the Galactic centre. Stars with $\langle L_{z}\rangle$ between 1 and 3 show stars that are probably well constrained within the MW’s disc, whereas those with 0 $\leq$ $\langle L_{z}\rangle$$<1$ have a higher probability of not being part of the disc, and those with $\langle L_{z}\rangle$$<0$ probably come from the MW’s halo. A uniform range is maintained on both the $x$ and $y$-axes, serving to direct the reader’s attention towards outliers manifesting within lower metallicities. The colour scheme remains consistent with the preceding figures.
@@ 3.1.2 Assessing thin and thick disc membership
Table 2: Table expressing the number of stars classified in the thin or thick disc according to the prescription adopted here for each metallicity-stratified group (‘G’).
@@ 3.1.3 The final sample post chemo-dynamic assessment
It is noteworthy that, after excluding stars potentially belonging to the Galactic halo and those that seem to belong to the thick disc, the median age, $\overline{t}_{\star}$, for the groups changes, and the variation becomes smaller. The difference between the youngest and oldest groups is reduced to $\sim$ 2 Gyr, which is within the expected age uncertainty ($\approx$ 1-2 Gyr; see Mints & Hekker, 2018). These are median values, and we discuss other effects of age and kinematics in Sect. 3.3. Additionally, Group 4, which is the second most metal-poor, is now the oldest when comparing their age distributions. We display both the Kiel diagram for entire final sample and the distribution of these ages in Figs. 10 and 11, respectively. The tracks shown in Fig. 10 are PARSEC isochrones retrieved from the CMD web interface. [^9] http://stev.oapd.inaf.it/cgi-bin/cmd Regarding Fig. 11, we note that this distribution is slightly different from the one in the adjacent upper axis of Fig. 2, which contains the full sample, not only the stars in the Galactic thin disc.
@@
Figure 11: Age distribution for the stars belonging to the Galactic disc illustrated in the shape of Gaussian kernel densities for easier view (instead of histograms). The $x$-axis has been truncated to represent a physically meaningful range of stellar ages. The vertical dashed lines in corresponding colours show the computed median ages for each subgroup in our disc stellar sample. The colour scheme is consistent with the preceding figures.
3.2 Birth radii estimation versus current Galactocentric distances
@@
Figure 12: Estimated Galactocentric birth radii, $\langle R_{\rm{b}}\rangle$, in contrast to the present guiding radii, $\langle R_{\rm{g}}\rangle$. Both radii are measured in kiloparsecs and encompass all stars in our sample that are part of the disc, according to Fig. 7. Each distinct subplot delineates a specific star group, classified using hierarchical clustering (as illustrated in Fig. 1). The subplots are arranged in descending order of the median iron abundance, $\langle\rm{[Fe/H]}\rangle$. At the upper part of each subplot, $\langle\rm{[Fe/H]}\rangle$ and $\overline{t}_{\star}$ are displayed. Furthermore, the overall standard deviations ($\sigma$) for $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$ are exhibited at the lower right corner of each subplot. We overlay the 2D-Gaussian kernel densities to illustrate the concentration patterns of the stars in each subplot. A reference black dashed line is incorporated, signifying the point of parity between both radii ($\langle R_{\rm{b}}\rangle$=$\langle R_{\rm{g}}\rangle$). The colour scheme remains consistent with the preceding figures.
@@
3.2.1 Challenges in predicting birth radii: Limitations of the chemical enrichment models and the generalised additive model
Upon examining the groups depicted in Fig. 12, particularly those with sub-solar metallicities (i.e. groups 5, 4, and 3), we observe that some stars exhibit relatively large birth radii ($\langle R_{\rm{b}}\rangle$ $\sim$ 28 kpc), exceeding the estimated radius of the MW, which is approximately 25 kpc (as previously discussed in Sect. 2.3.3). One possible explanation is contamination from thick disc stars, which poses challenges for elimination without significantly affecting the completeness of the thin disc population. However, if these stars predominantly belong to the thin disc, as seems more likely for the majority, alternative explanations may apply. Rather than interpreting this as evidence of stars originating outside the Galactic disc, it more likely reflects limitations in the chemical evolution models’ ability to accurately account for the age-metallicity relationship in certain stars. Estimating accurate stellar ages, particularly for field stars, remains a significant challenge. We must also consider potential mismatches between model predictions and the actual metallicity evolution in the outer regions of the Galactic thin disc. These combined limitations lead to predictions that some stars have birth radii larger than the known extent of the MW. While the precise birth radii for these stars may be inaccurate, the broader conclusion – that they formed in the outer disc and migrated inward – remains robust.
@@ 3.2.2 Galactocentric migration patterns across different stellar groups
Figure 12 shows that $\overline{t}_{\star}$ decreases as $\langle\rm{[Fe/H]}\rangle$ increases (except for Group 3), accompanied by a reduction in the overall variance of both $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$. In other words, as we look at groups with larger metallicities, the stars seem to cluster and become more tightly grouped within a narrower range of both $\langle R_{\rm{b}}\rangle$ and $\langle R_{\rm{g}}\rangle$ (we note in particular that while the range of $\langle R_{\rm{g}}\rangle$ values is the same in all panels, the range of $\langle R_{\rm{b}}\rangle$ values increases from the top left to the bottom right). The clustering in $\langle R_{\rm{b}}\rangle$ for the stars with high metallicity can be easily understood with a glance at the gradients displayed in Figs. 5 and 6. Old stars of high metallicity can only form in the inner regions of the Galaxy (as we discussed previously in Dantas et al., 2023). In addition, we note that a decrease in the number density of metal-rich stars towards the outer disc can be seen in the sample of other surveys as well (e.g. Hayden et al., 2015; Gaia Collaboration et al., 2023).
@@ 3.2.3 Other signs of migration
Figure 13: Variation of Galactocentric distance versus median eccentricity ($\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$  $\times$ $\langle e\rangle$) across stellar groups, arranged in descending order of median metallicity $\langle\rm{[Fe/H]}\rangle$. The range of eccentricity ($\langle e\rangle$) on the $x$-axis remains consistent across all subplots, while the $y$-axis retains an unconstrained scale, enhancing the visualisation of the pronounced $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$ dispersion. Corresponding to Fig. 12, a horizontal dashed line is positioned at $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$=0 to accentuate the region where Galactocentric distances coincide. To illuminate the 2D distribution of stars within the designated parameter space, 2D-Gaussian kernel densities are overlaid upon the scatter markers. Consistent with prior visuals, the colours in this figure maintain the same scheme.
Figure 14: Median maximum Galactic azimuthal height, $\langle Z_{\rm{max}}\rangle$, versus the median eccentricity, $\langle e\rangle$, for all disc stars in the six groups, ordered by decreasing metallicity from top left to bottom right. The 1D-Gaussian kernel density distributions for $\langle Z_{\rm{max}}\rangle$ and $\langle e\rangle$ are shown adjacent to the main subplots, with medians marked by dotted lines and annotated values. The main subplots display all disc stars, with thin (300 pc) and thick disc (900 pc) scale heights (McMillan, 2017) marked by dashed and dot-dashed lines, respectively, and a 2D Gaussian kernel density plot indicating higher density regions. Consistent $x$- and $y$-axis ranges highlight the shift in distributions across metallicity groups, with the proportion of stars with larger $\langle Z_{\rm{max}}\rangle$ increasing as metallicity decreases.
@@
Therefore, we analyse our stars’ $\langle e\rangle$ and $\langle Z_{\rm{max}}\rangle$ to explore whether these parameters are connected to migration patterns. Figure 13 presents the variations in radii, $\langle R_{\rm{g}}\rangle$-$\langle R_{\rm{b}}\rangle$ (or simply, $\Delta R$), against $\langle e\rangle$ for all groups in our sample. This figure provides alternative an view of the results shown in Fig. 12 by incorporating $\langle e\rangle$. In Fig. 13, the eccentricities of stars across all groups appear constrained to $\sim$0.4, though Group 6 contains a few outliers with higher eccentricities. As we move towards metal-poor groups, the variation in $\Delta R$ increases, but $\langle e\rangle$ remains largely unchanged.
@@
Another way to visualise the behaviour of the stars in our sample is by plotting $\langle Z_{\rm{max}}\rangle$ versus $\langle e\rangle$, which is displayed in Fig. 14. The subplots in Fig. 14 are organised by decreasing metallicity, from top left to bottom right. The scale heights of the thin and thick discs are indicated in the figure by dashed and dot-dashed lines at 0.3 and 0.9 kpc, respectively (estimates from McMillan, 2017). The consistent axis ranges across all subplots allowed us to observe the shifting distributions as metallicity decreases, notably the increasing proportion of stars with larger $\langle Z_{\rm{max}}\rangle$ values as metallicity decreases, especially for Groups 4 and 3. In fact, the median $\langle Z_{\rm{max}}\rangle$ for Group 4 is approximately at the scale height of the thick disc, and significantly above it for Group 5. The standard deviations for both $\langle e\rangle$ and $\langle Z_{\rm{max}}\rangle$ are displayed in each subplot, demonstrating that the variance in $\langle Z_{\rm{max}}\rangle$ increases as metallicity decreases, while the variance in $\langle e\rangle$ remains pretty constant. As for $\langle e\rangle$, it is noticeable that the medians are all $<0.2$, suggesting that most stars have near-circular orbits, but all have somewhat low $\langle e\rangle$, almost all having $\langle e\rangle$ $<0.4$; few outliers can be seen in Groups 2 and 6.
@@
As metallicity decreases, a greater number of stars appear to reach larger $\langle Z_{\rm{max}}\rangle$ values. This could be attributed to residual interlopers from the thick disc not previously identified in Sect. 3.1.2. However, if most of these stars indeed belong to the Galactic thin disc, as expected, this trend could be potentially explained by two factors. (i) Stars in the most metal-poor groups tend to concentrate at older ages (as is discussed in Sect. 3.3.1 and detailed in Table 10), giving them more time to interact with Galactic structures (e.g. giant molecular clouds), which may contribute to disc thickening, as shown by Roškar et al. (2013). However, this applies mainly to stars churned outward, and as we discuss in Sect. 3.3, the proportion of such stars in these metal-poor groups is low. (ii) Alternatively, this trend can be due to the interaction between the outer disc and the first pericentric passage of Sagittarius, which occurred around 6 Gyr ago. According to Das et al. (2024), stars older than 6 Gyr are significantly kinematically hotter than their younger counterparts. Indeed, our observations indicate that the stars in the most metal-poor groups – those likely formed at greater Galactocentric distances – have higher $\langle Z_{\rm{max}}\rangle$ values. This is consistent with the hypothesis that stars formed at larger radii are more susceptible to the influence of interactions with Sagittarius, leading to dynamic heating. Furthermore, as we show and discuss in Sect. 3.3, a substantial proportion of metal-poor stars appear to be migrating inward from the outskirts of the MW.
@@
To further explore this hypothesis, we depict $\langle Z_{\rm{max}}\rangle$ versus $\overline{t}_{\star}$ in Fig. 15, stratified by the HC groups. It is evident in Fig. 15 that stars aged 6 Gyr or older exhibit a higher proportion of larger $\langle Z_{\rm{max}}\rangle$ values, especially in the more metal-poor groups, such as Groups 5, 4, and 3. Overall, if the majority of these stars are not interlopers from the thick disc, we propose that this increased variation in $\langle Z_{\rm{max}}\rangle$ is likely due to a combination of outward churning and the dynamic influence of the pericentric passage of Sagittarius.